# Operational research approaches and mathematical models for kidney exchange: A literature survey and empirical evaluation

### v1

@@ 4.2.1 Results of Standard Implementations of the Combined Models
In Tables 3, 4 and 5 we present the results of our first set of experiments. In all three tables, the rows correspond to the cycle models, and the columns correspond to the chain models (for conciseness, we omit the affixes -CYCLE and -CHAIN). Moreover, in the rows called “AVG”, we present averages taken over all cycle models for each chain model, and similarly, in the columns called “AVG”, we present averages taken over all chain models for all cycle models. In addition, in the final row of each table we present the results for EF-HYBRID. In Table 3, we present for each combined model the number of instances solved to optimality within the time limit (column “#opt”) and the average CPU time in seconds across all instances including the ones were the time limit was hit (column “t”). Whenever the memory limit is reached, the instance is counted as unsolved and a time of 3600 seconds is used to compute the average CPU time. In Table 4 we present for each combined model the average number of variables (column “#v”) and constraints (column “#c”) in thousands. Whenever constraint generation is used (as indicated in Table 8 in Appendix A.4), we only count the constraints that are actually added to the model. For this table we only consider the 1251 unweighted instances where all combined models could be built (but not necessarily solved) in the time and memory limit. In Table 5 we present for each combined model the average absolute gap between the optimal solution value and the optimal value of the model’s LP relaxation for the unweighted instances and for the weighted instances. Note that for the models where constraint generation is used, the LP relaxation is solved before any additional constraints are added to the model. For this table we only consider the 2098 instances where the LP relaxation of each combined model could be solved in the time and memory limit for both the unweighted and weighted variants of the instance.
@@
Overall, the most effective combined models are HCF-CYCLE+PIEF-CHAIN and PIEF-CYCLE+PIEF-CHAIN, which both solved almost 85% of the tested instances to optimality. Furthermore, we found that the effectiveness of a cycle model is largely independent of the specific chain model it is combined with, and vice versa. Therefore, we proceed by focusing primarily on evaluating the cycle and chain models independently. It is also clear that the choice of chain model has a bigger impact on performance than the choice of cycle model. The full results show that this is particularly the case when $L$ is large with respect to $K$.
@@
Comparing the different cycle models, HCF-CYCLE and PIEF-CYCLE perform best overall, followed by CF-CYCLE and EEF-CYCLE, and the least effective model is clearly EF-CYCLE. We observe that CF-CYCLE, HCF-CYCLE and PIEF-CYCLE have the strongest LP relaxation. In fact, Dickerson et al., (2016) and Delorme et al., (2023) proved that the LP relaxations of these models are equally tight when not considering chains. EEF-CYCLE closely follows in terms of LP relaxation, and EF-CYCLE is far behind [^8] This is partly due to our implementation of constraint generation., likely explaining the relatively poor performance of the latter model, despite it having the smallest number of variables across all cycle models. Furthermore, CF-CYCLE has the largest number of variables across all cycle models, but this is partly compensated by the fact that it also has the smallest number of constraints. Conversely, HCF-CYCLE has, by far, the largest number of constraints but a more moderate number of variables. PIEF-CYCLE, in comparison, strikes a balance with a moderate number of both variables and constraints. Moreover, in contrast to what could be expected based on the worst-case upper bounds on the number of variables and constraints as presented in Table 8 in Appendix A.4, the model size of PIEF-CYCLE is smaller than that of EEF-CYCLE. This difference could be attributed to the ordering of vertices: EEF-CYCLE orders vertices in ascending order of degree to benefit its LP relaxation, while PIEF-CYCLE employs a descending order to minimise its model size.
@@
When examining the chain models, PIEF-CHAIN stands out as the best-performing model by a significant margin, consistently outperforming all other chain models. HCF-CHAIN and EF-CHAIN-MTZ rank as the second and third most effective chain models, respectively. To understand these results, note that PIEF-CHAIN has the strongest LP relaxation along with CF-CHAIN and HCF-CHAIN, while the LP relaxations of the EEF-based chain models are weaker and those of the EF-based chain models are considerably weaker. Our results are in line with a theoretical result from Dickerson et al., (2016), stating that the LP relaxation of CF-CYCLE+CF-CHAIN dominates that of CF-CYCLE+PIEF-CHAIN. However, in practice, the difference is only marginal. Despite its strong LP relaxation, CF-CHAIN suffers from having the largest number of variables across all chain models by far, which is also not compensated by the fact that CF-CHAIN is the chain model with the smallest number of constraints. The number of variables in HCF-CHAIN is substantially smaller than that of CF-CHAIN, while the number of constraints in the former model is only slightly larger than that of the latter, which explains why HCF-CHAIN outperforms CF-CHAIN by a clear margin. Furthermore, note that EF-CHAIN-MTZ is substantially more effective than EF-CHAIN-EXP, whereas the difference between the performance of EEF-CHAIN-EXP and EEF-CHAIN-MTZ is less pronounced. Interestingly, EF-CHAIN-EXP performs better than both EEF-based chain models. The main reason seems to be that the EEF-based chain models suffer from a very large number of variables compared to the other chain models, whereas the EF-based chain models have the smallest number of variables overall. The number of variables in PIEF-CHAIN is only slightly larger than that of the EF-based chain models on average, while its number of constraints is also among the smallest across all chain models, which along with its strong LP relaxation, explains the superior performance of PIEF-CHAIN.
@@
4.2.2 Results of RCVF Implementations
In Table 6, we present the performance of the most relevant model combinations when solved using RCVF. Specifically, we report the number of instances solved to optimality within the time limit (column “#opt”) and the average CPU time across all instances (column “t”), in both the unweighted and the weighted case. To provide a comparison, the numbers in brackets indicate the performance of the standard implementations of these models. In addition, we present for each model the average number of iterations required by RCVF (measured as the number of calls to the ILP solver, column “#iter”) and the percentage of variables remaining in the reduced ILP model at the final RCVF iteration relative to the full model (column “%vrem”). To compute “#iter” and “%vrem” we only considered the 2298 instances where all models could be solved to optimality for both the unweighted and weighted variant of the instance.
Table 6: Performance of the RCVF implementations across 2160 unweighted instances and 2160 weighted instances.
unweighted instances weighted instances
model #opt t #iter %vrem #opt t #iter %vrem
One key reason why the weighted instances are harder to solve than the unweighted instances, especially for RCVF, is that the LP relaxations of all models are significantly less tight in the weighted case (see Table 5). As a result, RCVF requires more iterations for weighted instances than for unweighted ones. Relating to this, the number of iterations is mostly consistent across all models for a given instance, with the PIEF-CHAIN models sometimes requiring a few more iterations compared to the CF-CHAIN models due to the LP relaxations of the former models being marginally weaker than that of the latter. On the other hand, the reduction in the number of variables due to RCVF is substantially larger for the weighted case compared to the unweighted case. The reduction is also larger for the CF-CHAIN models compared to the PIEF-CHAIN models, in particular for weighted instances, explaining why the CF-CHAIN models still benefit from RCVF in that case. Finally, while we only report statistics on the number of iterations and percentage of remaining variables for solved instances, we note that the unweighted instances where the time limit is reached are characterised by a percentage of remaining variables that is typically larger than that of solved unweighted instances. Conversely, the weighted instances where the time limit is reached are characterised by a number of iterations that is generally larger compared to that of solved weighted instances.
@@ 4.2.3 Results of Third-Party Methods
Next, in Table 7 we present the performance of the tested third-party methods. Namely, for each method we report again the number of instances solved to optimality within the time limit (column “#opt”) and the average CPU time across all instances (column “t”).
Our philosophy in evaluating the third-party methods was to use the codes as provided, modifying them as little as possible. Unfortunately, for some of the third-party methods, we experienced crashes (for reasons other than hitting the memory limit), that—to the best of our knowledge—were unrelated to our experimental setup. In each case, the code raised and captured an error. Moreover, to ensure a fair evaluation, we implemented a series of checks to validate the correctness and consistency of the reported solutions. Specifically, we verified that the returned objective values respected the best known lower bounds and upper bounds, that the reported time was less than the time limit whenever a solution was labelled as optimal, and conversely that the reported time was (approximately) equal to the time limit when no optimal solution was returned. While all our methods passed these checks, we did find that, for some of the instances, this was not the case for one or more of the tested third-party methods. Therefore, we quantified the number of instances where such inconsistencies occurred in columns “inconsistencies”, where we also indicate the type of inconsistency. Specifically, “obj $<$ LBbest” denotes that the objective value returned by a method is lower than the highest known lower bound, while “obj $>$ UBbest” denotes that the objective value returned by a method is higher than the lowest known upper bound. Whenever an inconsistency occurred, the instance is counted as unsolved and a time of 3600 seconds is used to compute the average CPU time.
Table 7: Performance of the third-party methods across 2160 unweighted instances and 2160 weighted instances.
@@
BNP-DFS 1067 1911 6$\times$ “obj $>$ UBbest 1068 1894 8$\times$ “obj $>$ UBbest”, 6$\times$ “obj $<$ LBbest
BNP-PICEF 1906 611 - 1320 1458 20$\times$ “obj $<$ LBbest
@@
JL-BNP 2152 24 4$\times$ “crashed” 1376 1375 299$\times$ “crashed”
JL-BNP-PICEF 2132 182 10$\times$ “crashed” 1291 1506 352$\times$ “crashed”, 14$\times$ “obj $<$ LBbest
*) Recall that DCD only applies to unweighted instances with $L=K$, and that it does not allow for chains of length 1.
$\dagger$) We chose to still treat these 14 instances as solved, since the returned solutions were all less than 0.01% away from optimal.
In the unweighted case, JL-BNP and JL-BNP-PICEF stand out as the most effective methods, both solving almost all of the instances to optimality. Built upon CF-CYCLE+CF-CHAIN and CF-CYCLE+PIEF-CHAIN, respectively, these methods improve substantially upon the standard implementation and the RCVF implementation of these combined models, indicating that B&P is a very powerful tool in this setting. We note that JL-BNP and JL-BNP-PICEF include many of the ideas that were proposed in the papers that introduced BNP-DFS, DCD, BP-MDD and BNP-PICEF. Furthermore, recall that JL-BNP and JL-BNP-PICEF use the solver Gurobi, whereas the other third-party methods use CPLEX.
Conversely, none of the third-party methods are competitive with the most effective standard implementations of the combined models in the weighted case. Among the third-party methods, JL-BNP remains the most effective, solving about 64% of the weighted instances to optimality, compared to 82% for the standard implementation of PIEF-CYCLE+PIEF-CHAIN. However, both JL-BNP and JL-BNP-PICEF suffer from a high number of crashes, which, in particular, causes JL-BNP-PICEF to solve slightly fewer instances than BNP-PICEF.
@@ 4.3 Summary of Results
In the experimental part of the survey, we emphasised that one can model the cycles-and-chains case of KE-Opt by combining any cycle model with any chain model. We experimentally evaluated all 35 model combinations that arose from combining any of the 5 cycle models with any of the 7 chain models that were discussed in Section 3. We also considered a hybrid model, RCVF enhancements and third-party implementations. Our main findings (including results presented in Appendix A.8) are as follows:
@@ 5 Conclusions and Future Directions
  • ILP models for KE-Opt. From an OR perspective, modelling and solving KE-Opt to enable optimal sets of exchanges to be found efficiently (in relation to both time and space) is a key challenge. Through increased participation in KEPs, and growth in infrastructure for conducting exchanges, we can expect larger pools, and longer cycles and chains, to feature going forward. These advances will ensure that designing algorithms to construct optimal solutions to KE-Opt will continue to be an important research direction. Although much progress has been made, as described in Sections 2.2 and 3, there is still scope for new ideas and techniques to emerge. Moreover, it remains open to further explore bespoke ILP models for specific KEP applications, such as in international KEPs where there may be additional country-specific constraints (Mincu et al., , 2021), or in the presence of specific hierarchical objectives (Delorme et al., , 2024).
  • Approximation and $\mathcal{FPT}$ algorithms. Although most KEPs focus on finding solutions that are optimal (for example in relation to the number of transplants or the total weight of the solution) at a given matching run, it is not always obvious that this is the best strategy in the long term Carvalho et al., (2024). Moreover as detailed in the previous item, KE-Opt is likely to become more challenging over time, with increased instance sizes, and feasible solutions allowing longer cycles and chains. The difficulty of finding optimal solutions may give further motivation for approximation algorithms for KE-Opt with good worst-case performance guarantees, where there is still a large gap between the best known upper and lower bounds as detailed in Section 2.1, and especially where these algorithms can be shown to outperform these worst-case guarantees in simulations. Moreover, perhaps more from a theoretical standpoint, there is still scope to further explore the parameterised complexity landscape for KE-Opt, where there have only been a handful of results so far, as outlined in Section 2.1.
@@ Acknowledgements
The authors would like to thank John Dickerson and Xenia Klimentova for making code available to us for our empirical evaluation. The authors would also like to thank Kristína Gáliková, Daniel Paulusma, Tuomas Sandholm and Danielius Šukys for helpful discussions regarding this paper. Rachael Colley, Maxence Delorme, David Manlove and William Pettersson were supported by grant EP/X013618/1 from the Engineering and Physical Sciences Research Council. Part of the work on this paper was carried out whilst Mathijs Barkel visited the University of Glasgow, supported by a grant from the CentER Graduate School of Tilburg University. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
Appendix A Appendices
A.1 EEF-CHAIN-EXP and EEF-CHAIN-MTZ
@@
For both EEF-based chain models, we consider $|\mathcal{N}|$ subgraphs, where for every $s\in\mathcal{N}$, subgraph $\mathcal{G}^{s}=(\mathcal{V}^{s},\mathcal{A}^{s})$ is the graph induced by $\mathcal{V}^{s}=\{s\}\cup\mathcal{R}\cup\{\tau\}$. Subsequently, we introduce a binary decision variable $y_{uv}^{s}$ for every arc $(u,v)\in\mathcal{A}^{s}$ in every subgraph $s\in\mathcal{N}$, taking value $1$ if arc $(u,v)$ is selected in subgraph $s$, and value 00 otherwise. This allows us to efficiently forbid chains of length exceeding $L$ by limiting the number of selected arcs in each subgraph to at most $L$. However, as in the case of EF-CHAIN-EXP and EF-CHAIN-MTZ, we must also exclude all cycles. To that end, we present again two models that exclude cycles in a different way.
In the first model, EEF-CHAIN-EXP, we forbid cycles using an exponential number of constraints. Note that this is in contrast to EEF-CYCLE, which is fully polynomial. Defining $\mathcal{C}_{\leq L-2}$ as the set of cycles in $\mathcal{G}$ (that is, the original compatibility graph) of length at most $L-2$, we obtain the following:
[…]
@@
$\sum_{(u,v)\in\mathcal{A}^{s}}y_{uv}^{s}\leq L\cdot\sum_{v:(s,v)\in\mathcal{A}^{s}\setminus\{(s,\tau)\}}y_{sv}^{s}$ $\forall s\in\mathcal{N},$ (49)
@@
The objective function (45) maximises the total weight and constraints (46) and (47) enforce that all RDPs and NDDs, respectively, are involved in at most one chain. Moreover, constraints (48) ensure flow conservation and constraints (49) limit the number of selected arcs per subgraph to at most $L$, which guarantees that the maximum chain length is never exceeded. These constraints also break symmetry by ensuring that if any arc is selected in some subgraph, then an arc leaving the NDD in that subgraph (but not going to $\tau$) must be selected as well. However, constraints (46)-(49) and (51) do not forbid the selection of multiple exchanges per subgraph, similar to what could happen in EEF-CYCLE. In this case, for every subgraph it is still allowed to select one or more cycles whose lengths add up to at most $L-2$. Therefore, constraints (50) are added to forbid all such cycles. The EEF-based chain model by Anderson et al., (2015) contains different cycle-elimination constraints, which we comment on in Appendix A.2.
@@
[…]
$t_{u}-t_{v}+(L-1)\sum_{\begin{subarray}{c}s\in\mathcal{N}:\\ (u,v)\in\mathcal{A}^{s}\end{subarray}}y^{s}_{uv}+(L-3)\sum_{\begin{subarray}{c}s\in\mathcal{N}:\\ (v,u)\in\mathcal{A}^{s}\end{subarray}}y^{s}_{vu}\leq L-2$
$\forall(u,v)\in\mathcal{A}_{\mathcal{R}},$ (53)
$1\displaystyle 1\leq t_{v}\leq L-$ 1 $\forall v\in\mathcal{R}.$ (54)
Constraints (A.1) and (54) simultaneously exclude chains of length more than $L$ and cycles of any length.
Example 1.
(continued) Consider the instance presented in Figure 3 for $L=4$. The subgraphs $\mathcal{G}^{5}$ and $\mathcal{G}^{6}$ required by EEF-CHAIN-EXP and EEF-CHAIN-MTZ are presented in Figure 7. The total number of arcs across both subgraphs is $24$, leading to $24$ variables for EEF-CHAIN-EXP. EEF-CHAIN-MTZ has $4$ additional timestamp variables, namely one per RDP. For instance, chain $\left\langle 5,1,2,3,\tau\right\rangle$ is obtained by setting $y_{51}^{5}=y_{12}^{5}=y_{23}^{5}=y_{3\tau}=1$ and for EEF-CHAIN-MTZ also $t_{1}=1$, $t_{2}=2$, $t_{3}=3$ and $t_{4}\in[1,3]$. Note that chain $\left\langle 6,2,3,4,1,\tau\right\rangle$ is infeasible due to constraints (49), as this chain would require too many arcs to be selected in subgraph $\mathcal{G}^{6}$. Moreover, cycle $\left\langle 1,4,1\right\rangle$ is excluded directly for EEF-CHAIN-EXP due to the constraint (50) corresponding to this cycle, while it is excluded indirectly for EEF-CHAIN-MTZ via the timestamp variables.
[…]
We found that it is not computationally advantageous (or expedient) to disaggregate constraints (50) or (A.1) (i.e., to add a constraint per subgraph) as this would both increase the number of constraints in the model and weaken the LP-relaxation. Furthermore, constraint generation may be required to deal with constraints (50), as there is an exponential number of them.
A.2 Alternative Constraints for the EF- and EEF-based Models

### v2

@@ 4.2.1 Results of Standard Implementations of the Combined Models
In Tables 3, 4 and 5 we present the results of our first set of experiments. In all three tables, the rows correspond to the cycle models, and the columns correspond to the chain models (for conciseness, we omit the affixes -CYCLE and -CHAIN). Moreover, in the rows called “AVG”, we present averages taken over all cycle models for each chain model, and similarly, in the columns called “AVG”, we present averages taken over all chain models for each cycle model. In addition, in the final row of each table we present the results for EF-HYBRID. In Table 3, we present for each combined model the number of instances solved to optimality within the time limit (column “#opt”) and the average CPU time in seconds across all instances including the ones were the time limit was hit (column “t”). Whenever the memory limit is reached, the instance is counted as unsolved and a time of 3600 seconds is used to compute the average CPU time. In Table 4 we present for each combined model the average number of variables (column “#v”) and constraints (column “#c”) in thousands. Whenever constraint generation is used (as indicated in Table 7 in Appendix E), we only count the constraints that are actually added to the model. For this table we only consider the 1251 unweighted instances where all combined models could be built (but not necessarily solved) in the time and memory limit. In Table 5 we present for each combined model the average absolute gap between the optimal solution value and the optimal value of the model’s LP relaxation for the unweighted instances and for the weighted instances. Note that for the models where constraint generation is used, the LP relaxation is solved before any additional constraints are added to the model. For this table we only consider the 2098 instances where the LP relaxation of each combined model could be solved in the time and memory limit for both the unweighted and weighted variant of the instance. Given the highly heterogeneous nature of the instances considered, we emphasise that the main purpose of these tables is to compare the relative performance of multiple models, rather than to assess the performance of an individual model.
@@
*) To put these gaps into perspective, the average optimal value is 103.65 for the unweighted instances and 6431.65 for the weighted instances.
@@
Overall, the most effective combined models are HCF-CYCLE+PIEF-CHAIN and PIEF-CYCLE+PIEF-CHAIN, which both solved almost 85% of the tested instances to optimality. Furthermore, we found that the relative effectiveness of a cycle model compared to other cycle models is largely independent of the specific chain model that is used, and vice versa. Therefore, we proceed by focusing primarily on evaluating the cycle and chain models independently. It is also clear that the choice of chain model has a bigger impact on performance than the choice of cycle model. The full results show that this is particularly the case when $L$ is large with respect to $K$.
@@
Comparing the different cycle models, HCF-CYCLE and PIEF-CYCLE perform best overall, followed by CF-CYCLE and EEF-CYCLE, and the least effective model is clearly EF-CYCLE. As could be expected from the results in Section 3.8, model combinations involving CF-CYCLE, HCF-CYCLE and PIEF-CYCLE have the strongest LP relaxation, while combinations involving EEF-CYCLE closely follow in terms of LP relaxation, and combinations with EF-CYCLE are far behind [^13] This is partly because we report the value of the LP relaxation before adding any additional constraints to the model., likely explaining the relatively poor performance of the latter cycle model, despite it having the smallest number of variables across all cycle models. Furthermore, CF-CYCLE has the largest number of variables across all cycle models, but this is partly compensated by the fact that it also has the smallest number of constraints. Conversely, HCF-CYCLE has, by far, the largest number of constraints but a more moderate number of variables. PIEF-CYCLE, in comparison, strikes a balance with a moderate number of both variables and constraints. Moreover, in contrast to what could be expected based on the worst-case upper bounds on the number of variables and constraints as presented in Table 7 in Appendix E, the model size of PIEF-CYCLE is smaller than that of EEF-CYCLE. This difference could be attributed to the ordering of vertices: EEF-CYCLE orders vertices in ascending order of degree to benefit its LP relaxation, while PIEF-CYCLE employs a descending order to minimise its model size.
@@
When examining the chain models, PIEF-CHAIN stands out as the best-performing model by a significant margin, consistently outperforming all other chain models. HCF-CHAIN and EF-CHAIN-MTZ rank as the second and third most effective chain models, respectively. To understand these results, note that model combinations involving PIEF-CHAIN have the strongest LP relaxation along with combinations involving CF-CHAIN and HCF-CHAIN, while the LP relaxations of combinations involving EEF-based chain models are weaker and those involving the EF-based chain models are considerably weaker. In line with the observations by Dickerson et al., (2016), we observe that the LP relaxation of CF-CYCLE+CF-CHAIN is in practice only marginally stronger than that of CF-CYCLE+PIEF-CHAIN (see also Section 3.8). Despite its strong LP relaxation, CF-CHAIN suffers from having the largest number of variables across all chain models by far, which is also not compensated by the fact that CF-CHAIN is the chain model with the smallest number of constraints. The number of variables in HCF-CHAIN is substantially smaller than that of CF-CHAIN, while the number of constraints in the former model is only slightly larger than that of the latter, which explains why HCF-CHAIN outperforms CF-CHAIN by a clear margin. Furthermore, note that EF-CHAIN-MTZ is substantially more effective than EF-CHAIN-EXP, whereas the difference between the performance of EEF-CHAIN-EXP and EEF-CHAIN-MTZ is less pronounced. Interestingly, EF-CHAIN-EXP performs better than both EEF-based chain models. The main reason seems to be that the EEF-based chain models suffer from a very large number of variables compared to the other chain models, whereas the EF-based chain models have the smallest number of variables overall. The number of variables in PIEF-CHAIN is only slightly larger than that of the EF-based chain models on average, while its number of constraints is also among the smallest across all chain models, which along with its strong LP relaxation, explains the superior performance of PIEF-CHAIN.
@@ 4.2.2 Results of Third-Party Methods
Next, in Table 6 we present the performance of the third-party methods indicated in Table 2. Namely, for each method we report again the number of instances solved to optimality within the time limit (column “#opt”) and the average CPU time across all instances (column “t”).
Our philosophy in evaluating the third-party methods was to use the codes as provided, modifying them as little as possible. Unfortunately, for some of the third-party methods, we experienced crashes (for reasons other than hitting the memory limit), that—to the best of our knowledge—were unrelated to our experimental setup. In each case, the code raised and captured an error. Moreover, to ensure a fair evaluation, we implemented a series of checks to validate the correctness and consistency of the reported solutions. Specifically, we verified that the returned objective values respected the best known lower bounds and upper bounds, that the reported time was less than the time limit whenever a solution was labelled as optimal, and conversely that the reported time was (approximately) equal to the time limit when no optimal solution was returned. While all our methods passed these checks, we did find that, for some of the instances, this was not the case for one or more of the tested third-party methods. Therefore, we quantified the number of instances where such inconsistencies occurred in columns “inconsistencies”, where we also indicate the type of inconsistency. Specifically, “obj << LBbest” denotes that the objective value returned by a method is lower than the highest known lower bound, while “obj >> UBbest” denotes that the objective value returned by a method is higher than the lowest known upper bound. Whenever an inconsistency occurred, the instance is counted as unsolved and a time of 3600 seconds is used to compute the average CPU time.
Table 6: Performance of the third-party methods across 2160 unweighted and 2160 weighted instances.
@@
[…]
@@
[…]
*) Recall that DCD only applies to unweighted instances with $L=K$, and that it does not allow for chains of length 1. $\dagger$) We chose to still treat these 14 instances as solved, since the returned solutions were all less than 0.01% away from optimal.
In the unweighted case, JL-BNP and JL-BNP-PICEF stand out as the most effective methods, both solving almost all of the instances to optimality. Built upon CF-CYCLE+CF-CHAIN and CF-CYCLE+PIEF-CHAIN, respectively, these methods improve substantially upon the standard implementation and the RCVF implementation (see Appendix I) of these combined models, indicating that B&P is a very powerful tool in this setting. We note that JL-BNP and JL-BNP-PICEF include many of the ideas that were proposed in the papers that introduced BNP-DFS, DCD, BP-MDD and BNP-PICEF. Furthermore, recall that JL-BNP and JL-BNP-PICEF use the solver Gurobi, whereas the other third-party methods use CPLEX.
Conversely, none of the third-party methods are competitive with the most effective standard implementations of the combined models in the weighted case. Among the third-party methods, JL-BNP remains the most effective, solving about 64% of the weighted instances to optimality, compared to 82% for the standard implementation of PIEF-CYCLE+PIEF-CHAIN. One explanation for this difference in performance is the high number of crashes for JL-BNP and JL-BNP-PICEF. In particular, we find that while JL-BNP and JL-BNP-PICEF excel on instances where the LP relaxation is tight (i.e., instances where the optimal solution value matches the dual bound obtained after solving the root node relaxation), they underperform on instances where the LP relaxation is weaker—a feature that characterises weighted instances more so than unweighted ones (see Table 5).
@@ 4.3 Summary of Results
In the experimental part of the survey, we emphasised that one can model the cycles-and-chains case of KE-Opt by combining any cycle model with any chain model. We experimentally evaluated all 35 model combinations that arose from combining any of the 5 cycle models with any of the 7 chain models that were discussed in Section 3. We also considered a hybrid model, RCVF enhancements and third-party implementations. Our main findings (including results presented in Appendices I and J) are as follows:
@@ 5 Conclusions and Future Directions
  • Dealing with challenging KE-Opt instances. From an OR perspective, modelling and solving KE-Opt to enable optimal sets of exchanges to be found efficiently (in relation to both time and space) is a key challenge. Through increased participation in KEPs, and growth in infrastructure for conducting exchanges, we can expect larger pools, and longer cycles and chains, to feature going forward. These advances will mean that designing algorithms to construct optimal solutions to KE-Opt will continue to be an important research direction. Although much progress has been made, as described in Sections 2.2 and 3, there is still scope for new ideas and techniques to emerge. Moreover, it remains open to further explore bespoke ILP models for specific KEP applications, such as in international KEPs where there may be additional country-specific constraints (Mincu et al., , 2021), or in the presence of specific hierarchical objectives (Delorme et al., , 2024). The challenging nature of KE-Opt instances may give further motivation for approximation algorithms for KE-Opt with good worst-case performance guarantees, where there is still a large gap between the best known upper and lower bounds as detailed in Section 2.1, and especially where these algorithms can be shown to outperform these worst-case guarantees in simulations. Further, perhaps more from a theoretical standpoint, there is still scope to further explore the parameterised complexity landscape for KE-Opt, where there have only been a handful of results so far, as outlined in Section 2.1.
@@ Acknowledgements
The authors would like to thank the Editor-in-Chief and three Anonymous Reviewers for their support of this work, and for their detailed and valuable comments that have helped to improve the presentation of this paper. We would also like to thank John Dickerson and Xenia Klimentova for making code available to us for our empirical evaluation. Further, we would like to thank Margarida Carvalho, Kristína Gáliková, Sushmita Gupta, Daniel Paulusma, Tuomas Sandholm, Danielius Šukys and Ana Viana for helpful suggestions and discussions regarding this paper. Rachael Colley, Maxence Delorme, David Manlove and William Pettersson were supported by grant EP/X013618/1 from the Engineering and Physical Sciences Research Council. Part of the work on this paper was carried out whilst Mathijs Barkel visited the University of Glasgow, supported by a grant from the CentER Graduate School of Tilburg University. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
Appendices
Appendix A The EEF-CHAIN-EXP and EEF-CHAIN-MTZ Models
@@
For both EEF-based chain models, we consider $|\mathcal{N}|$ subgraphs, where for every $s\in\mathcal{N}$, subgraph $\mathcal{G}^{s}=(\mathcal{V}^{s},\mathcal{A}^{s})$ is the graph induced by $\mathcal{V}^{s}=\{s\}\cup\mathcal{R}\cup\{\tau\}$. Subsequently, we introduce a binary decision variable $y_{uv}^{s}$ for every arc $(u,v)\in\mathcal{A}^{s}$ in every subgraph $s\in\mathcal{N}$, taking value $1$ if arc $(u,v)$ is selected in subgraph $s$, and value 0 otherwise. This allows us to efficiently forbid chains of length exceeding $L$ by limiting the number of selected arcs in each subgraph to at most $L$. However, as in the case of EF-CHAIN-EXP and EF-CHAIN-MTZ, we must also exclude all cycles. To that end, we present again two models that exclude cycles in a different way.
In the first model, EEF-CHAIN-EXP, we forbid cycles using an exponential number of constraints. Note that this is in contrast to EEF-CYCLE, which is fully polynomial. Defining $\mathcal{C}_{\leq L-2}$ as the set of cycles in $\mathcal{G}$ (that is, the original compatibility graph) of length at most L2L-2, we obtain the following:
$(\texttt{EEF-CHAIN-EXP})\quad$ $\max$ $\sum_{s\in\mathcal{N}}\sum_{(u,v)\in\mathcal{A}^{s}}w_{uv}y_{uv}^{s}$ (45)
@@
$\sum_{(u,v)\in\mathcal{A}^{s}}y_{uv}^{s}\leq L\cdot\sum_{v:(s,v)\in\mathcal{A}^{s}\setminus\{(s,\tau)\}}\hskip-15.0pty_{sv}^{s}$ $\forall s\in\mathcal{N},$ (49)
@@
The objective function (45) maximises the total weight and constraints (46) and (47) enforce , respectively, that each NDD and each RDP is involved in at most one chain. Moreover, constraints (48) ensure flow conservation and constraints (49) limit the number of selected arcs per subgraph to at most $L$, which guarantees that the maximum chain length is never exceeded. These constraints also break symmetry by ensuring that if any arc is selected in some subgraph, then an arc leaving the NDD in that subgraph (but not going to $\tau$) must be selected as well. However, constraints (46)-(49) and (51) do not forbid the selection of multiple exchanges per subgraph, similar to what could happen in EEF-CYCLE. In this case, for every subgraph it is still allowed to select one or more cycles whose lengths add up to at most L2L-2. Therefore, constraints (50) are added to forbid all such cycles. The EEF-based chain model by Anderson et al., 2015b contains different cycle-elimination constraints, which we comment on in Appendix B.
@@
$t_{u}-t_{v}+(L-1)\hskip-6.0pt\sum_{\begin{subarray}{c}s\in\mathcal{N}:\\ (u,v)\in\mathcal{A}^{s}\end{subarray}}\hskip-6.0pty^{s}_{uv}+(L-3)\hskip-6.0pt\sum_{\begin{subarray}{c}s\in\mathcal{N}:\\ (v,u)\in\mathcal{A}^{s}\end{subarray}}\hskip-6.0pty^{s}_{vu}\leq L-2$
$\hskip 199.16928pt\forall(u,v)\in\mathcal{A}_{\mathcal{R}},$ (53)
$1\leq t_{v}\leq L-1\hskip 129.46011pt\forall v\in\mathcal{R}.$ (54)
Constraints (53) and (54) simultaneously exclude chains of length more than $L$ and cycles of any length.
We found that it is not computationally advantageous (or expedient) to disaggregate constraints (50) or (53) (i.e., to add a constraint per subgraph) as this would both increase the number of constraints in the model and weaken the LP-relaxation. Furthermore, constraint generation may be required to deal with constraints (50), as there is an exponential number of them.
Appendix B Alternative Constraints for the EF- and EEF-based Models